OpenStudy (anonymous):

express -2rad(2) + 2rad(2)i in polar form

OpenStudy (anonymous):

\[a+ib\] For the complex numbers: Magnitude is given by: \[M = \sqrt{(a^2 + b^2)}\] And Arguement is given by: \[\theta = \tan^{−1}\frac{b}{a}\]

OpenStudy (anonymous):

so the magnitude is 4

OpenStudy (anonymous):

and the argument is -24.9356.....how does that help?

OpenStudy (anonymous):

Magnitude is 4.. No you need to solve into decimal just keep it in fraction..

OpenStudy (anonymous):

i dont understand the part about solving into decimal